OpenStudy (anonymous):

can someone solve this? heping my friend on her homework. best answer gets medal 5x^-3y^2/x^5y^-1 * (2xy^3)^-2/xy

OpenStudy (anonymous):

do you mean \[(5x^{3}-3y^{2}/x^{5}y^{-1})\times((2xy^{3})^{2}/xy) or (5x^{3}-3y^{2}/x^{5}y^{-1}\times(2xy^{3})^{2}/xy)\]

OpenStudy (anonymous):

\[5x ^{-3}y ^{2}/x ^{5}y ^{-1}\times(2xy ^{3})^{-2}/xy\]

OpenStudy (anonymous):

@DarthTony

OpenStudy (anonymous):

\[\large \frac{ 5x^{-3}y^{2}}{ x^5y^{-1} }\times \frac{ (2xy^3)^{-2} }{ xy }=5x^{(-3)-(5)}y^{(2)-(-1)}\times \frac{ 2x^{-2}y^{3(-2)} }{ xy }\] \[\large =5x^{-8}y^{2+1}\times \frac{2x^{-2}y^{-6}}{xy}\] \[\large =5x^{-8}y^3 \times 2x^{(-2)-(1)}y^{(-6)-(1)}\] \[\large =5x^{-8}y^3 \times 2x^{-3}y^{-7}\] \[\large =(5\times 2)x^{(-8)+(-3)}\times y^{(3)+(-7)}\]

OpenStudy (anonymous):

\[\large =10x^{-11}\times y^{3-7}\] \[\large =10x^{-11}\times y^{-4}\] \[\large =10x^{-11}y^{-4}\]